# Random-Set Neural Networks (RS-NN)

### v2

@@ D.5.4 Credal set width
D.5.5 Ablation study on hyperparameters
As we discussed in the main paper, the regularization serves the purpose of ensuring that our model generates valid belief functions, which are constrained by maintaining a sum of masses equal to 1 and ensuring non-negativity. These terms are designed to penalize deviations from valid belief functions. However, it is crucial not to assign too much weight to this term, as excessively penalizing deviations may hinder the model’s ability to accurately classify data points. For example, in a Variational Auto Encoder (VAE), if we assign too much weight to the KL divergence term in the loss, the model may prioritize fitting the latent distribution at the expense of reconstructing the input data accurately. This imbalance can lead to poor reconstruction quality and suboptimal performance on downstream tasks.
@@ D.5.5 Ablation study on hyperparameters
In Fig. 23, we show the test accuracies for different values of hyperparameters $\alpha$ and $\beta$ in $\mathcal{L}_{B-RS}$ loss function (7). Hyperparameters $\alpha$ and $\beta$ adjusts the relative significance of the regularization terms $M_{r}$ and $M_{s}$ respectively in $\mathcal{L}_{B-RS}$ loss. The test accuracies are calculated for CIFAR-10 dataset with varying $\alpha$ (blue) and $\beta$ (red) values, $\alpha/\beta$ = [0.001, 0.005, 0.006, 0.009, 0.01, 0.015, 0.020, 0.025, 0.03, 0.04, 0.05]. Test accuracy is the highest when $\alpha$ and $\beta$ equals $0.5$.
@@
In our experiments across various datasets and architectures, we found that a value of 1e-3 for the hyperparameters yields satisfactory results. This includes architectures ranging from ResNet-50 to Vision Transformers (ViT-Base), and datasets ranging from MNIST ($\approx$ 60,000 images of 10 classes) to ImageNet ($\approx$ 1.1M images of 1000 classes). While conducting a parameter search for each dataset could potentially lead to further optimization, it’s worth noting that, in many cases, this step can be omitted without sacrificing performance significantly. For further optimization one can, of course, perform a tailored parameter search per dataset, but this is not quite necessary.
@@
Figure 23: Test accuracies of RS-NN on different values of hyperparameters $\alpha$ and $\beta$ in the loss function.
D.5.6 Ablation Study on the number of focal sets
@@
Figure 24: Ablation study on number of non-singleton focal sets $K$ on CIFAR-10 dataset using ViT-Base. The maximum value of K can be 1013 for 10 classes (after excluding the singletons and empty set).
@@
The number of non-singleton focal sets $K$ to be budgeted is a hyperparameter and needs to be studied. A smaller value of $K$ can lead to more similar results to classical classification while a larger value of $K$ can increase the complexity. Therefore, we conducted an ablation study on the $K$ on the CIFAR-10 dataset. It was found that a small value of $K$ (comparable to the number of classes) works best and even performs better than $K=0$ (when there are no non-singleton focal sets) (see Fig. 24). Hence, using set prediction along with the proposed budgeting algorithm not only helps induce uncertainty quantification but also improves the performance of the model. Note that performance comparison was made in terms of accuracy. Smet’s Pignistic Transform smets05ijar was used to compute class-wise probabilities from the predicted belief function. Fig. 24 shows an ablation study on different number of focal sets $K$ for CIFAR-10 on the Vision-Transformer (ViT-Base) model.

### v3

@@ E.5.2 Credal set width
E.5.3 Entropy vs Credal Set Width
There has been considerable research on uncertainty measures for credal sets. One of the earlier works by Yager (2008) makes the distinction between two types of uncertainty within a credal set: conflict (also known as randomness or discord) and non-specificity. Non-specificity essentially varies with the size of the credal set (Kolmogorov, 1965). Since by definition, aleatoric uncertainty refers to the inherent randomness or variability in the data, while epistemic uncertainty relates to the lack of knowledge or information about the system (Hüllermeier & Waegeman, 2021), conflict and non-specificity directly parallel these concepts. These measures of uncertainty are axiomatically justified (Bronevich & Klir, 2008).
Figs. 24(a) and 24(b) depict the relationship between the entropy of RS-NN pignistic prediction and the associated confidence level (Fig. 24(a)), and between credal set width and confidence (Fig. 24(b)) for the CIFAR-10 (left) and SVHN/Intel Image (right) datasets. The distribution of iD predictions (left) for both tests show a concentration at the top left, indicating high confidence and low entropy or credal set width. Conversely, OoD predictions (middle, right) exhibit a more dispersed pattern.
Entropy, reflecting prediction uncertainty, is quite correlated with confidence, in both iD and OoD tests. In contrast, as it considers the entire set of plausible outcomes within a belief function rather than a single prediction, credal set width better quantifies the degree of epistemic uncertainty inherent to a prediction. As a result, credal set width is less dependent on the concentration of predictions and is more reflective of the overall uncertainty encompassed by the model. Fig. 24(b) shows that, unlike entropy, credal with is clearly not correlated with confidence.
E.6 Ablation study on hyperparameters $\alpha$ and $\beta$
The regularization serves the purpose of ensuring that our model generates valid belief functions, which are constrained by maintaining a sum of masses equal to 1 and ensuring non-negativity. These terms are designed to penalize deviations from valid belief functions. However, it is crucial not to assign too much weight to this term, as excessively penalizing deviations may hinder the model’s ability to accurately classify data points. For example, in a Variational Auto Encoder (VAE), if we assign too much weight to the KL divergence term in the loss, the model may prioritize fitting the latent distribution at the expense of reconstructing the input data accurately. This imbalance can lead to poor reconstruction quality and suboptimal performance on downstream tasks.
@@ E.6 Ablation study on hyperparameters $\alpha$ and $\beta$
In Fig. 25, we show the test accuracies for different values of hyperparameters $\alpha$ and $\beta$ in $\mathcal{L}_{B-RS}$ loss function (Eq. 7). Hyperparameters $\alpha$ and $\beta$ adjusts the relative significance of the regularization terms $M_{r}$ and $M_{s}$ respectively in $\mathcal{L}_{B-RS}$ loss. The test accuracies are calculated for CIFAR-10 dataset with a fixed number of focal sets $K=20$ and varying $\alpha$ (blue) and $\beta$ (red) values, $\alpha/\beta$ = [0.001, 0.005, 0.006, 0.009, 0.01, 0.015, 0.020, 0.025, 0.03, 0.04, 0.05]. Test accuracy is the highest when $\alpha$ and $\beta$ equals $31e-$ 3.
@@
In our experiments across various datasets and architectures, we found that a value of $31e-$ 3 for the hyperparameters yields satisfactory results. This includes architectures ranging from ResNet-50 to Vision Transformers (ViT-Base-16), and datasets ranging from MNIST ($\approx$ 60,000 images of 10 classes) to ImageNet ($\approx$ 1.1M images of 1000 classes). While conducting a parameter search for each dataset could potentially lead to further optimization, it is worth noting that, in many cases, this step can be omitted without sacrificing performance significantly. For further optimization one can, of course, perform a tailored parameter search per dataset, but this is not quite necessary.
@@
Figure 25: Test accuracies of RS-NN on the CIFAR-10 dataset using ResNet50, for a fixed $K=20$ and different values of hyperparameters $\alpha$ and $\beta$ in the loss function.
An alternative loss with valid belief functions. A straightforward way of enforcing the positivity of the masses is to apply softmax over the masses computed in the BCE loss. This ensures that the masses are non-negative and sum to one. After obtaining the softmaxed masses, one can compute the belief functions from these normalized masses and minimize the loss between this computed belief and the original unnormalized belief.
The random set loss $\mathcal{L}_{RS}$ now becomes,
$\mathcal{L}_{RS}=\mathcal{L}_{BCE}+\mathcal{L}_{BCE_{norm}},$
where $\mathcal{L}_{BCE}$ is the BCE loss between belief function logits and ground truth, and $\mathcal{L}_{BCE_{norm}}$ is the BCE loss between the normalized belief function logits reconstructed from mass logits that have been softmaxed.
We implemented this simple method on ResNet50 RS-NN on the CIFAR-10 dataset and obtained an accuracy of $92.47\%$. This is quite close to our original accuracy of $93.53\%$ with the mass regularizations. This mirrors results in the literature showing that soft constraints work as well as hard ones in practice (Márquez-Neila et al., 2017).
E.7 Ablation Study on the number of focal sets
@@
Figure 26: Ablation study on number of non-singleton focal sets $K$ on CIFAR-10 dataset using ViT-Base-16 (with $\alpha$ = $\beta$ = $31e-$ 3). The maximum value of K can be 1013 for 10 classes (after excluding the singletons and empty set).
@@
The number of non-singleton focal sets $K$ to be budgeted is a hyperparameter and needs to be studied. A smaller value of $K$ can lead to more similar results to classical classification while a larger value of $K$ can increase the complexity. Therefore, we conducted an ablation study on $K$ on the CIFAR-10 dataset. We found that a small value of $K$ (comparable to the number of classes) works best and even performs better than $K=0$ (when there are no non-singleton focal sets) (see Fig. 26). Hence, using set prediction along with the proposed budgeting algorithm not only helps induce uncertainty quantification but also improves the performance of the model. Note that performance comparison was made in terms of accuracy. Smet’s Pignistic Transform (Smets, 2005) was used to compute class-wise probabilities from the predicted belief function. Fig. 26 shows an ablation study on different number of focal sets $K$ for CIFAR-10 on the Vision-Transformer (ViT-Base-16) model with $\alpha=\beta=1e-3$.
To further support our claim, we compare a budgeted RS-NN with standard RS-NN (full $2^{N}$ sets) on the CIFAR-10 datasets using ResNet50 as the backbone. In Tab. 16 below, we report the test accuracy, OoD detection metrics (AUROC, AUPRC), pignistic entropy and credal set width for the standard RS-NN vs budgeted RS-NN. The standard RS-NN has 1024 ($2^{10}$) sets since CIFAR-10 has 10 classes, and the budgeted RS-NN has $K=20$ focal (non-singleton) sets, so $10+20=30$ input sets. Tab. 16 shows that budgeted RS-NN performs better than the standard model, especially iD vs OoD entropy, and AUROC and AUPRC scores.
As we say in the main paper, while, in theory, using the full power set is ideal for uncertainty estimation, in practice, having all those degrees of freedom might make it more difficult for the network to learn. When a model is confronted with an overwhelming number of input sets, it may struggle to identify meaningful patterns amidst the noise created by less relevant or redundant combinations. This excessive flexibility can hinder the network’s ability to converge effectively, as it may become trapped in local minima or overfit to the training data. This can dilute the model’s capacity to generalize well to unseen data, thereby complicating the learning of the underlying relationships between features and the set structure of the data.
Table 16: Comparison of accuracy, AUROC, AUPRC, pignistic entropy, and credal set width for standard RS-NN vs budgeted RS-NN on the CIFAR-10 dataset.
In-distribution (iD) Out-of-distribution (OoD)
Dataset Model Test accuracy (%) ($\uparrow$) ECE ($\downarrow$) SVHN Intel Image
AUROC ($\uparrow$) AUPRC ($\uparrow$) AUROC ($\uparrow$) AUPRC ($\uparrow$)
CIFAR-10 Standard RS-NN $92.66$ $0.0501$ $93.39$ $91.38$ $95.84$ $89.33$
Budgeted RS-NN $\mathbf{93.53}$ $\mathbf{0.0484}$ $\mathbf{94.91}$ $\mathbf{93.72}$ $\mathbf{97.39}$ $\mathbf{90.27}$
SVHN Intel Image
Entropy (iD) ($\downarrow$) Credal Set Width (iD) ($\downarrow$) Entropy ($\uparrow$) Credal Set Width ($\uparrow$) Entropy ($\uparrow$) Credal Set Width ($\uparrow$)
Standard RS-NN $0.286\pm 0.80$ $0.048\pm 0.15$ $\mathbf{1.205}\pm\mathbf{1.20}$ $\mathbf{0.408}\pm\mathbf{0.07}$ $1.490\pm 0.71$ $\mathbf{0.669}\pm\mathbf{0.21}$
Budgeted RS-NN $\mathbf{0.088}\pm\mathbf{0.308}$ $\mathbf{0.007}\pm\mathbf{0.044}$ $1.132\pm 0.855$ $0.260\pm 0.322$ $\mathbf{1.517}\pm\mathbf{0.740}$ $0.587\pm 0.367$
E.8 On the feasibility of budgeting of sets
Tabs. 17 and 18 show that replacing t-SNE with UMAP for budgeting in the RS-NN model yields comparable performance in terms of test accuracy, out-of-distribution (OoD) detection, and uncertainty estimation. Both t-SNE and UMAP are capable of embedding data in a 3D space, with UMAP employing a faster algorithm that utilizes five nearest neighbors and a minimum distance of 0.9. This approach ensures that the integrity of the data structure is maintained while significantly reducing computational costs. UMAP offers an efficient alternative without compromising the model’s effectiveness. To make the generation of embeddings for budgeting of focal sets more efficient, we can utilize faster alternatives like UMAP instead of t-SNE, without compromising the overall model performance.
Table 17: Test accuracy and OoD detection (AUROC, AUPRC) results on RS-NN where budgeting is done using t-SNE vs budgeting done on UMAP.
In-distribution (iD) Out-of-distribution (OoD)
Dataset Model Test accuracy (%) ($\uparrow$) ECE ($\downarrow$) SVHN Intel Image
AUROC ($\uparrow$) AUPRC ($\uparrow$) AUROC ($\uparrow$) AUPRC ($\uparrow$)
CIFAR-10 RS-NN (t-SNE) 93.53 0.0484 94.91 93.72 97.39 90.27
RS-NN (UMAP) 92.97 0.0482 94.31 92.46 96.22 89.66
Table 18: Entropy and credal set width results on RS-NN where budgeting is done using t-SNE vs budgeting done on UMAP. The goal is to minimize both metrics for in-distribution (iD) data while increasing them for out-of-distribution (OoD) data.
In-distribution (iD) Out-of-distribution (OoD)
Dataset Model Entropy ($\downarrow$) Credal Set Width ($\downarrow$) SVHN Intel Image
Entropy ($\uparrow$) Credal Set Width ($\uparrow$) Entropy ($\uparrow$) Credal Set Width ($\uparrow$)
CIFAR-10 RS-NN (t-SNE) 0.088 $\pm$ 0.308 0.007 $\pm$ 0.044 1.132 $\pm$ 0.855 0.260 $\pm$ 0.322 1.517 $\pm$ 0.740 0.587 $\pm$ 0.367
RS-NN (UMAP) 0.082 $\pm$ 0.305 0.007 $\pm$ 0.047 1.119 $\pm$ 0.887 0.341 $\pm$ 0.374 1.423 $\pm$ 0.762 0.582 $\pm$ 0.402
A 2D visualization of the ellipses formed by calculating the principle axes and their lengths using eigenvectors and eigenvalues obtained from GMMs, for each class $c$ (Spruyt, 2013) is shown in Fig. 27.
The budgeting procedure can also be modified to accommodate multiple data sources or a continuous stream of data. t-SNE can be replaced by an autoencoder (this will need to be trained first) or PCA to add continual inference capabilities in the dimensionality reduction step. Similarly, Gaussian-based dynamic probabilistic clustering (GDPC) (Diaz-Rozo et al., 2018) can be used instead of GMM clustering to handle a continuous stream of data. GDPC works by first initialising a GMM and then updating its parameters as it sees new data. In alternative, (Kulis & Jordan, 2011) have proposed an interesting approach to learn a GMM from multiple sources of data by maintaining a local and a global mixture model. Both these approaches can be plugged into our budgeting framework with little modifications to add continual learning capabilities to it. The computation of cluster overlaps remains the same, so overlapping scores and the resultant focal sets will need to be updated as the clusters evolve. However, a cluster tracking strategy (Barbara & Chen, 2001) can be employed so that the overlap assessment step is only re-done when a sufficient drift has been detected in the clusters. All the proposed changes are efficient enough to not have a significant effect on the overall time of budgeting.
Appendix F Application on text classification
The random-set approach can be applied on any classification task. In this section, we detail experiments performed on text classification.
Dataset. We chose the BBC text dataset (Greene & Cunningham, 2006), which consists of various news categories such as tech, business, sport, and entertainment, and used this data for training our model. The task is to classify the given text into one of these categories. The dataset is structured with two columns: category and text, where the category is the label, and the text is the content to be classified. The text data will be preprocessed and fed into the model to predict the category of each text.
Model. To train our model, we leveraged a pre-trained BERT model available on TensorFlow Hub. Specifically, we used the small BERT ($L-4_{H}-512_{A}-8$) model, which is a lightweight version of BERT. The model is designed to take text input, preprocess it with BERT’s preprocessing layer, pass it through BERT’s encoder to generate embeddings, and then use a dropout layer. The final layer is a fully-connected layer with sigmoid activation for the RS-NN BERT classifier (RS-NN BERT), and a fully-connected layer with softmax activation for the standard BERT classifier (CNN BERT).
Training. We trained these models on the BBC text dataset, fine-tuning the BERT model as part of the training process. Both models were trained for 10 epochs using Adam optimizer, a batch size of 32, and a learning rate of 3e-5. RS-NN BERT has all the sets of classes excluding the null set and full set ($2^{5}$ = 32 - 2 = 30 classes) and CNN BERT has the 5 original classes.
Out-of-distribution detection. For OoD detection, we use the Emotion Detection from Text (Seyeditabari et al., 2018) dataset which contains tweets annotated with emotional labels. The dataset includes three columns: tweetid, sentiment, and content, where sentiment represents the emotion behind each tweet. The dataset includes 13 emotion classes such as anger, fear, joy, love, sadness, and surprise, etc, aiming to identify emotional expressions in text.
Experimental results. In Tab. 19 below, we show the test accuracy, out-of-distribution (OoD) metrics (AUROC, AUPRC), and the in-distribution (iD) vs OoD entropy for RS-NN BERT and CNN BERT. RS-NN achieves higher overall performance with a significantly higher AUROC and AUPRC scores, highlighting the efficiency of the model at differentiating between iD and OoD samples. RS-NN has higher OoD entropy than CNN, but also has higher iD entropy than CNN, which indicates that the uncertainty in predictions is higher as it is a small dataset.
Table 19: Test accuracy, AUROC, AUPRC, iD vs OoD entropy for RS-NN BERT and CNN BERT (both fine-tuned on BERT).
Dataset (iD) Model Test Acc. (%) ($\uparrow$) iD Entropy ($\downarrow$) Emotion Detection (OoD)
AUROC ($\uparrow$) AUPRC ($\uparrow$) Entropy ($\uparrow$)
BBC Text (iD) RS-NN BERT 96.85 0.541 $\pm$ 0.196 95.19 95.93 1.510 $\pm$ 0.611
CNN BERT 94.15 0.027 $\pm$ 0.125 56.71 57.54 0.059 $\pm$ 0.191